2x-2y=-12
4x-7y=-15

Answer by checkley75(3666)

MULTIPLY EQUATION 1 BY -2 (-4X+4Y=24) NOW ADD THESE 2 EQUATIONS
-4X+4Y=24
4X-7Y=-15
----------------------------
-3Y=9
Y=9/-3
Y=-3 ANSWER.
4X-7*-3=-15
4X+21=-15
4X=-15-21
4X=-36
X=-36/4
X=-9 ANSWER.
PROOF
2*-9-2*-3=-12
-18+6=-12
-12=-12